Listed below are the salaries, in thousands of dollars, for a sample of 15 chief financial officers in the electronics industry: $554.0 577.0 498.0 540.0 509.0 582.0 534.0 497.0 523.0 581.0 559.0 506.0 493.0 543.0 499.0 a. Determine the mean, the median, and the standard deviation. (Round the final answers to 2 decimal places.) Mean Median Standard deviation b. Determine the coefficient of skewness. (Round the final answer to 3 decimal places.) Coefficient of skewness
Answer #1

= (554+ 577+498+.....+499) /15

= $ 533

number of observation = 15

median is the 8th oobservation after arranging the data in ascending order

median = $534

=[(554-533)2 +.....+(499-533)2]/14=  1042.143

Standard deviation = square root of variance = $32.28

coefficient of skewness = (mean -median )/ Standard deviation = -0.031
Answer #1

= (554+ 577+498+.....+499) /15

= $ 533

number of observation = 15

median is the 8th oobservation after arranging the data in ascending order

median = $534

=[(554-533)2 +.....+(499-533)2]/14=  1042.143

Standard deviation = square root of variance = $32.28

coefficient of skewness = (mean -median )/ Standard deviation = -0.031